Fin d th e chang e i n th e interna l energ y o f on e mol e o f a 

monatomic ideal gas in an isobaric expansion at 1 atm from a volume 

of 5m 3

 t o a volume of 10m3

. y for a monatomi c idea l gas is 5/3 . 


1
Expert's answer
2021-04-05T11:12:14-0400

"\Delta U=\frac 32 p\Delta V,"

"\Delta U=\frac32\cdot 101325\cdot 5\approx 760~kJ."